iosangel:

1.) Line a passes through points (-2, 1) and (2, 9). Write an equation in slope intercept form that is parallel to line a.

mxddi3:

For parallel lines, the slopes must be equal, so the slope of the new line must also be the slope of line a. to do so, we must first find the slop of line a. this can be done by using the slope formula of \[\frac{ y2-y1 }{ x2-x1}\] so what is \[\frac{ 9-1 }{ 2- [-2] }\] ??

iosangel:

2

Shadow:

Point-Slope Form \[y - y_{1} = m(x - x_{1})\]

mxddi3:

We can plug the new slope and the given point into the slope-intercept form to solve for the y-intercept of the new line. so slope intercept form is \[y=mx+b\] we use ONE ordered pair, so i am just going to use the first one of (-2,1) 1=2(-2)+b 1=-4+b so then how do you solve the equation 1= -4+b ?

Shadow:

You want to use point slope form, not slope intercept.

Shadow:

You insert the slope, 2, in for m, then either ordered pair, and it will return a line that has the same slope (2x) but at a different y intercept.

Shadow:

Oh I see what you mean, that's one way to do it.

Shadow:

I just learned it the other way but that works fine.

mxddi3:

yes, i was gonna do A first so then they know what it shouldn't be and then we were gonna do your way lol

iosangel:

b is 5

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 yh, u can also do what Shadow said and use the point slope form: \[y-y1=m(x-x1) \] just use your slope of 2 and use an ordered pair. \(\color{#0cbb34}{\text{End of Quote}}\)

mxddi3:

^the other ordered pair

iosangel:

so 1=2(-2)+b could i write that as a answer

mxddi3:

well it says in slope intercept form and from what i've learned it would be y=mx+b (ex y=5x-2) so i think u have to write it like tht^